OpenStudy (anonymous):

What is the formula for factoring a^2+b^2?

OpenStudy (anonymous):

Specifically, my question is 8r^6+27s^12.

OpenStudy (anonymous):

It concerns factoring.

OpenStudy (anonymous):

\[a^2+b^2=a^2-(-b^2)=\left( a^2-\iota^2b^2 \right)=\left( a+\iota b \right)\left( a-\iota b \right)\]

OpenStudy (anonymous):

Is there a more simple version of this formula @surjithayer ?

OpenStudy (anonymous):

\[8r^6+27s^{12}=\left( 2r^2 \right)^3+\left( 3s^4 \right)^3=?\]

OpenStudy (anonymous):

Is that really the formula I need to use to solve the problem @surjithayer ?

OpenStudy (anonymous):

yes \[a^3+b^3=\left( a+b \right)\left( a^2-ab+b^2 \right)\]

OpenStudy (anonymous):

Ah, okay. Thank you. I am now going to work on it. :-)

OpenStudy (anonymous):

That was a hard question. (2r^2+3s^4)(4r^4-6r^4s^6+9s^3)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

R u asking what that is and its A²+b ²=c² and its called the pythagorean theorem